OpenStudy (anonymous):

The difference of two numbers is 5, and their product is 66. To find these numbers, let x equal the smaller. In terms of x, what is the larger number (their difference is five)?

myininaya (myininaya):

x-y=5 xy=66 assume x<y do you know how I got those two equations above?

OpenStudy (anonymous):

or just guess

OpenStudy (anonymous):

yes I do :)

OpenStudy (anonymous):

5 and 10 don't work because produce is 50

myininaya (myininaya):

ok so if x-y=5, does x=y+5?

OpenStudy (anonymous):

how about 6 and 11?

myininaya (myininaya):

shhh satellite